0.05 mL of 0.04020m sodium salicylate was diluted with iron 3 chloride solution in a 100ml volumetric flask. what is the conc of the iron 3 salicylate in the flask?
1
Expert's answer
2020-09-17T07:18:57-0400

3C7H5NaO3 + FeCl3 = (C7H5O3)3Fe + 3NaCl

"\rho" (C7H5NaO3) ≈ 1 g/ml

n (C7H5NaO3) = Cm x m(solvent) = 0.04020 x (0.05/1000)=0.00000201 mol

n ((C7H5O3)3Fe) = 1/3 x n (C7H5NaO3) = 0.00000201/3 = 0.00000067 mol

"\rho" ((C7H5O3)3Fe) ≈ 1 g/ml

Cm ((C7H5O3)3Fe) = 0.00000067 / ((100+0.05)/1000) = 0.000007 m